# On-device Sora: Enabling Diffusion-Based Text-to-Video Generation for Mobile Devices

### v1

@@ 4. Linear Proportional Leap
We reduce the excessive number of denoising steps performed by STDiT (Zheng et al., 2024) by introducing Linear Proportional Leap (LPL), which leverages the trajectory characteristics of Rectified Flow (Liu et al., 2022). It allows early stop of denoising steps through proportionally scaled linear leaps, without any extra model training or modification of the STDiT architecture.
4.1. Rectified Flow
@@
The reverse diffusion process is performed through multiple denoising steps, transforming an initial Gaussian distribution into a desired distribution corresponding to the input prompt. Several ODE-based methods (Zheng et al., 2023; Lu et al., 2022) reformulate this process by training neural network models to predict the drift at each time point, building the distribution trajectories from the initial to the target point by using ODE solvers (Runge, 1895; Yang et al., 2023).
@@
Rectified Flow (Liu et al., 2022) simplifies the transition from the initial point to the target point by training the model to predict a drift aligned with the direct linear trajectory connecting these two points. Using the Euler method (Chen et al., 2018), the $k$th trajectory $\boldsymbol{z_{k}}$ is derived by updating the previous trajectory $\boldsymbol{z_{k-1}}$ with the estimated drift $\boldsymbol{v}(P_{k},t_{k})$ and step size $dt_{k}$ defined by two sampled time steps $t_{k}$ and $t_{k+1}$, as follows:
(3) $\begin{split}\boldsymbol{z}_{k}&=\boldsymbol{z}_{k-1}+\boldsymbol{v}(P_{k},t_{k})dt_{k}~{}\forall 1\leq k\leq K\\ &~{}~{}\text{where}~{}t_{k}\in[0,1]~{}\text{and}~{}dt_{k}=\begin{cases}t_{k}-t_{k+1},&\text{if }t_{k}\neq t_{K}\\ t_{k},&\text{if }t_{k}=t_{K}.\end{cases}\end{split}$
@@
Here, the time step $t_{k}\in[0,1]$ corresponds to the normalized reverse process at the $k$th denoising step, with $t_{k}=1$ representing the time step at which data is fully noisy (start of denoising) and $t_{k}=0$ corresponding to the time step when data reaches the desired distribution (end of denoising). The drift $\boldsymbol{v}(P_{k},t_{k})$ is predicted from STDiT (Zheng et al., 2024) given the $k$th position on the trajectory, $P_{k}=t_{k}P_{1}+(1-t_{k})P_{0}$, computed from linear interpolation with sampled time step, $t_{k}$ (Liu et al., 2022).
@@
In Rectified Flow (Liu et al., 2022), the model is trained to predict the direct direction toward target point at any point on the trajectory. This allows diffusion-based image generation models for achieving a denoising process in few steps without significant performance degradation. While the actual trajectory may not be fully captured by a single drift, it can be effectively approximated with a reduced number of drifts. However, in diffusion-based video generation, achieving single-step generation is challenging due to the additional temporal variance of video data and the increased complexity of the target distribution. With Rectified Flow (Liu et al., 2022), Open-Sora (Zheng et al., 2024) generates video outputs with $K=30$ or $50$ denoising steps, in contrast to image generation, which produces images in a single step.
4.2. Linear Proportional Leap
Building upon Rectified Flow (Liu et al., 2022), Linear Proportional Leap reduces denoising steps, as illustrated in Figure 4.
@@ 4.2. Linear Proportional Leap
Figure 4. An abstracted illustration of trajectories and latent visualizations for $K=30$ and $n=15$: (a) Rectified Flow (Liu et al., 2022) with full $k=30$ denoising steps, generating intact and complete data, (b) Rectified Flow (Liu et al., 2022) with $n+1=16$ denoising steps without applying Linear Proportional Leap, resulting in low-quality data generation from variance with high step sizes ($dt_{k}$), and (c) Linear Proportional Leap with $n+1=15+1$ denoising steps, producing data nearly equivalent to (a).
@@
If the $n$th data distribution is sufficiently close to the $K$th target distribution in the denoising process, the trajectories $\boldsymbol{z}_{n+1...K}$ would be approximately straight for the remaining time steps $t_{n+1...K}$, making the estimation of the drift $\boldsymbol{v}(P_{k},t_{k})dt_{k}$ in Equation 3 unnecessary for $k>n$. Consequently, $\boldsymbol{v}(P_{k},t_{k})dt_{k}$ is estimated only for $1\leq k\leq n$, allowing the denoising process to stop early at the $(n+1)$th step, rather than continuing to the full $K$th step. For the remaining time steps $t_{n+1...K}$, the trajectories $\boldsymbol{z}_{n+1...K}$ linearly leap towards the target data distribution, with the straight direction of $\boldsymbol{v}(P_{n+1},t_{n+1})$ and $dt_{n+1}$ scaled proportionally to $t_{n+1...K}$.
By assuming $k=K$, $\boldsymbol{z}_{k}$ is derived from Equation 3 as:
(4) $\begin{split}\boldsymbol{z}_{k}&=\boldsymbol{z}_{k-1}+\boldsymbol{v}(P_{k},t_{k})dt_{k}\\ &=\boldsymbol{z}_{0}+\sum_{i=1}^{k-1}\boldsymbol{v}(P_{i},t_{i})({t_{i}-t_{i+1}})+\boldsymbol{v}(P_{k},t_{k})t_{k}\end{split}$
@@
If the denoising process stops at the step $n$, the $n$th trajectory can be represented as $\boldsymbol{z}_{n}=\boldsymbol{z}_{0}+\sum_{i=1}^{n}\boldsymbol{v}(P_{i},t_{i})({t_{i}-t_{i+1}})$. Then, if we apply the identical drift $\boldsymbol{v}(P_{n+1},t_{n+1})dt_{n+1}$ to the remaining $n+1\leq i\leq k$ steps, Equation 4 becomes:
(5) $\begin{split}\boldsymbol{z}_{k}&=\boldsymbol{z}_{n}+\boldsymbol{v}(P_{n+1},t_{n+1})\sum_{i={n+1}}^{k-1}(t_{i}-t_{i+1})+\boldsymbol{v}(P_{n+1},t_{n+1})t_{k}\\ &=\boldsymbol{z}_{n}+\boldsymbol{v}(P_{n+1},t_{n+1})(t_{n+1}-t_{n+2}+\cdots+t_{k-1}-t_{k}+t_{k})\\ &=\boldsymbol{z}_{n}+\boldsymbol{v}(P_{n+1},t_{n+1})t_{n+1}\end{split}$
@@
Thus, the required denoising steps are reduced to $n+1$ out of the total $k$ steps, allowing STDiT to be executed only $n+1$ times, with the last $(n+1)$th trajectory applied to its time step $t_{n+1}$. Replacing $dt_{k}$ with $t_{n+1}$ in Equation 5, instead of computing difference between the sampled time steps, can invoke an identical effect under assumption that later steps tend to sustain their drift directions. It enables the immediate completion of the denoising process, as $t_{n+1}$ is equivalent to the remaining steps required to reach the end of denoising.
For instance, when $K=30$ and $n=15$, the denoising steps proceed for the initial $n=15$ steps, and at step $n=16$, the trajectory transitions by leaping towards the linear direction of the target distribution. Due to the straightened trajectory, whose straightness increases as the denoising process approaches its end, combining the current drift with the product of the remaining step size yields a result that closely approximates the outcome achievable through the full denoising process. This enables a reduction of the denoising process by half without requiring additional model training.
@@
Linear Proportional Leap can be dynamically applied by measuring the cosine similarity between two consecutive drifts $\boldsymbol{v}$ at runtime. When the cosine similarity appears that the current trajectory is sufficiently linear, a linear leap is made proportionally to the remaining steps. Figure 5 visualizes the cosine similarities between consecutive drifts, which stabilizes after a certain number of steps, suggesting that the trajectory toward the target data distribution is nearly linear. This enables fewer steps by utilizing the leap with the larger step size to efficiently progress to the desired direction.
@@
Figure 5. An example of cosine similarities between two consecutive drifts estimated from STDiT (Zheng et al., 2024), i.e., $\boldsymbol{v}(P_{n},t_{n})$ and $\boldsymbol{v}(P_{n-1},t_{n-1})$ for 30 (red) and 50 steps (blue).
5. Temporal Dimension Token Merging
@@
We streamline the computational complexity of the denoising process by introducing Temporal Dimension Token Merging (TDTM), which halves the size of tokens within STDiT (Zheng et al., 2024) along the temporal dimension, decreasing the computation of the self-attention quadratically and the cross-attention in half. Unlike existing token merging primarily apply to self-attentions over the spatial dimension, which exhibits suboptimal performance (Bolya et al., 2022, 2023; Feng and Zhang, 2023; Li et al., 2024a), Temporal Dimension Token Merging leverages the temporal aspect of video frames to reduce computations while ensuring the video quality.
5.1. Token Merging
STDiT (Zheng et al., 2024) consists of multiple attention layers, i.e., cross- and self-attention, of the linear and quadratic complexity, respectively. In video generation, these attentions extend across two dimensions, i.e., spatial and temporal dimension.
@@ 5.1. Token Merging
General model optimization techniques, e.g., pruning (Reed, 1993), quantization (Gray and Neuhoff, 1998), and distillation (Hinton, 2015), may reduce the STDiT’s computational complexity. However, they often necessitate model re-training (fine-tuning) or specialized hardware for implementation, and most importantly, the performance of video generation can hardly be preserved. In contrast, token merging (Bolya et al., 2023) reduces the size of tokens processed in attention layers, decreasing computational complexity without requiring model re-training or hardware-specific adaptations.
@@
In STDiT (Zheng et al., 2024), the attention bias influenced by input tokens in cross-attention layers leads to higher computational demands compared to self-attention layers. As a result, developing an effective token merging method for cross-attention is crucial in video generation. However, existing token merging (Bolya et al., 2022, 2023; Li et al., 2024a) applied to the cross-attention have shown suboptimal performance, and applying them to the self-attention in STDiT has observed video quality drops (Bolya et al., 2022, 2023).
5.2. Temporal Dimension Token Merging
@@
Figure 6. In attention layers of STDiT (Zheng et al., 2024), two consecutive tokens are merged along the temporal dimension and subsequently unmerged after processing, reducing the token size by half and the computational complexity up to a quarter.
@@
Figure 6 illustrates Temporal Dimension Token Merging. Based on the hypothesis that successive video frames exhibit similar values, two consecutive frames are merged over the temporal dimension by averaging, creating a single token without an overhead of calculating frame similarity. This reduces the size of tokens by half while preserving the essential temporal information. Consequently, it decreases the computation of self-attention by a factor of four, according to the self-attention’s quadratic complexity, $\mathcal{O}(n^{2})$. Similarly, it reduces the computation of cross-attention by half, based on the cross-attention’s linear complexity, $\mathcal{O}(nm)$. Then, the output token processed through attention modules replicates the dimensions for each frame, restoring them to their original size. Figure 7 depicts the token merging and unmerging.
Figure 7. An illustration of the token merging and unmerging process over the temporal dimension.
@@
Given the token $\boldsymbol{T}_{in}$ as an input for a self- or cross-attention layer with the dimension [$B$, $ST$, $C$], where $B$ denotes the batch size, $S$ is the number of pixel patches, $T$ is the number of frames, and $C$ is the channel dimension, the input token $\boldsymbol{T}_{in}$ is merged into $\boldsymbol{T}_{merged}$, using the index $i$, as:
(6) $\boldsymbol{T}_{merged}=TDTM_{merge}(\boldsymbol{T}_{in})$
(7) $TDTM_{merge}(\boldsymbol{T})[i]=\frac{1}{2}(\boldsymbol{T}[:,i,:]+\boldsymbol{T}[:,i+1,:])$
@@
From this, two adjacent tokens are averaged along the temporal dimension, producing the merged token $\boldsymbol{T}_{merged}$ of [$B$, $ST/2$, $C$], which is processed by the self- or cross-attention.
@@
(8) $\boldsymbol{T}_{unmerged}=TDTM_{unmerge}(\text{Attention}(\boldsymbol{T}_{merged}))$
(9) $TDTM_{unmerge}(\boldsymbol{T})[2i]=\boldsymbol{T}[:,i,:]$

### v2

@@ 4 Linear Proportional Leap
On-device Sora reduces the excessive number of denoising steps performed by STDiT [83] by introducing Linear Proportional Leap (LPL), which leverages the trajectory properties of Rectified Flow [38]. It allows early stop of denoising steps through proportionally scaled linear leaps, without extra model training or modification of STDiT architecture.
4.1 Rectified Flow
@@
The reverse diffusion process is performed through multiple denoising steps, transforming an initial Gaussian distribution into a desired distribution corresponding to the input prompt. Several ODE-based methods [82, 40] reformulate this process by training models to predict the drift at each time point, building the distribution trajectories from the initial to the target point by using ODE solvers [52, 73].
@@
Rectified Flow [38] simplifies the transition from the initial point to the target point by training the model to predict a drift aligned with the direct linear trajectory connecting these two points. Using the Euler method [10], the $k$th trajectory $\boldsymbol{z_{k}}$ is derived by updating the previous trajectory $\boldsymbol{z_{k-1}}$ with the estimated drift $\boldsymbol{v}(P_{k},t_{k})$ and step size $dt_{k}$ defined by two sampled time steps $t_{k}$ and $t_{k+1}$, as follows:
$\begin{split}&\boldsymbol{z}_{k}=\boldsymbol{z}_{k-1}+\boldsymbol{v}(P_{k},t_{k})dt_{k}~{}~{}\forall 1\leq k\leq K\\ &\text{where}~{}t_{k}\in[0,1]~{}\text{and}~{}dt_{k}=\begin{cases}t_{k}{-}t_{k+1},&\text{if }t_{k}\neq t_{K}\\ t_{k},&\text{if }t_{k}=t_{K}\end{cases}\end{split}$ (3)
@@
Here, the time step $t_{k}{\in}[0,1]$ corresponds to the normalized reverse process at the $k$th denoising step, with $t_{k}{=}1$ representing the time step at which data is fully noisy (start of denoising) and $t_{k}{=}0$ corresponding to the time step when data reaches the desired distribution (end of denoising). The drift $\boldsymbol{v}(P_{k},t_{k})$ is predicted from STDiT [83] given the $k$th position on the trajectory, $P_{k}=t_{k}P_{1}+(1-t_{k})P_{0}$, computed from linear interpolation with sampled time step, $t_{k}$ [38].
@@
In Rectified Flow [38], the model is trained to predict the direct direction toward target point at any point on the trajectory. This allows diffusion-based generation models for achieving a denoising process in few steps without significant performance degradation. With Rectified Flow [38], Open-Sora [83] generates video with $K=30$ or $50$ steps.
4.2 Linear Proportional Leap
@@
Figure 3: An abstracted illustration of trajectories and latent visualizations for $K=30$ and $n=15$: (a) Rectified Flow [38] with full $k=30$ denoising steps, generating intact and complete video data, (b) Rectified Flow [38] with $n+1=16$ denoising steps without applying Linear Proportional Leap, resulting in low-quality video data generation from variance with high step sizes ($dt_{k}$), and (c) Linear Proportional Leap with $n+1=15+1$ denoising steps, producing video data nearly equivalent to (a).
Based on Rectified Flow [38], the proposed Linear Proportional Leap reduces denoising steps, as illustrated in Fig. 3.
@@
If the $n$th data distribution is sufficiently close to the $K$th target distribution in the denoising process, the trajectories $\boldsymbol{z}_{n+1...K}$ would be approximately straight for the remaining time steps $t_{n+1...K}$, making the drift estimation $\boldsymbol{v}(P_{k},t_{k})dt_{k}$ in Eq. 3 unnecessary for $k{>}n$. Consequently, $\boldsymbol{v}(P_{k},t_{k})dt_{k}$ is estimated only for $1{\leq}k{\leq}n$, allowing the denoising process to stop early at the $(n+1)$th step, not to the full $K$th step. For the remaining steps $t_{n+1...K}$, the trajectories $\boldsymbol{z}_{n+1...K}$ linearly leap towards the target data distribution, with the straight direction of $\boldsymbol{v}(P_{n+1},t_{n+1})$ and $dt_{n+1}$ scaled proportionally to $t_{n+1...K}$.
By assuming $k=K$, $\boldsymbol{z}_{k}$ is derived from Eq. 3 as:
$\begin{split}\boldsymbol{z}_{k}&=\boldsymbol{z}_{k-1}+\boldsymbol{v}(P_{k},t_{k})dt_{k}\\ &=\boldsymbol{z}_{0}+\sum_{i=1}^{k-1}\boldsymbol{v}(P_{i},t_{i})({t_{i}-t_{i+1}})+\boldsymbol{v}(P_{k},t_{k})t_{k}\end{split}$ (4)
@@
If the denoising process stops at the step $n$, the $n$th trajectory is represented as $\boldsymbol{z}_{n}=\boldsymbol{z}_{0}+\sum_{i=1}^{n}\boldsymbol{v}(P_{i},t_{i})({t_{i}-t_{i+1}})$. Then, if we apply the identical drift $\boldsymbol{v}(P_{n+1},t_{n+1})dt_{n+1}$ to the remaining $n+1\leq i\leq k$ steps, Eq. 4 becomes:
$\begin{split}\boldsymbol{z}_{k}&{=}\boldsymbol{z}_{n}{+}\boldsymbol{v}(P_{n+1},t_{n+1})\sum_{i={n+1}}^{k-1}(t_{i}{-}t_{i+1}){+}\boldsymbol{v}(P_{n+1},t_{n+1})t_{k}\\ &{=}\boldsymbol{z}_{n}{+}\boldsymbol{v}(P_{n+1},t_{n+1})(t_{n+1}{-}t_{n+2}{+}\cdots{+}t_{k-1}{-}t_{k}{+}t_{k})\\ &{=}\boldsymbol{z}_{n}{+}\boldsymbol{v}(P_{n+1},t_{n+1})t_{n+1}\end{split}$ (5)
@@
Thus, the required denoising steps are reduced to $n+1$ out of the total $k$ steps, allowing STDiT to be executed only $n+1$ times, with the last $(n+1)$th trajectory applied to its time step $t_{n+1}$, which enables Linear Proportional Leap. Replacing $dt_{k}$ with $t_{n+1}$ in Eq. 5, instead of computing difference between the sampled time steps, can invoke an identical effect under assumption that later steps tend to sustain drift directions. It enables the immediate completion of the denoising process, as $t_{n+1}$ is equivalent to the remaining denoising steps required to reach the end of denoising.
@@
Linear Proportional Leap can be dynamically applied by measuring the cosine similarity between two consecutive drifts $\boldsymbol{v}$ at runtime. When the cosine similarity appears that the current trajectory is sufficiently linear, a linear leap is made proportionally to the remaining steps. Fig. 4 visualizes the cosine similarities between consecutive drifts, stabilizing after a certain number of steps, suggesting that the trajectory toward the target data distribution is nearly linear. This enables fewer steps by utilizing the leap with the larger step size to efficiently progress to the desired direction.
@@
Figure 4: An example of cosine similarities between two adjacent drifts estimated from STDiT [83], i.e., $\boldsymbol{v}(P_{n},t_{n})$ and $\boldsymbol{v}(P_{n-1},t_{n-1})$ for 30 (red) and 50 steps (blue).
5 Temporal Dimension Token Merging
@@
On-device Sora reduces the computational complexity of the denoising process by introducing Temporal Dimension Token Merging (TDTM), which halves the size of tokens in STDiT [83] over the temporal dimension, decreasing the computation of self-attention quadratically and cross-attention in half. Unlike existing token merging applying to self-attentions over the spatial dimension, which exhibits suboptimal performance [6, 7, 17, 34], Temporal Dimension Token Merging leverages the temporal aspect of video frames to reduce computation while ensuring video quality.
5.1 Token Merging
@@
STDiT [83] consists of multiple attention layers, i.e., cross- and self-attention, of the linear and quadratic complexity, respectively. In video generation, these attentions extend across two dimensions, i.e., spatial and temporal dimension. General model optimization techniques, e.g., pruning [49], quantization [20], and distillation [22], may reduce STDiT’s attention computation. However, they necessitate model training (fine-tuning) or specialized hardware for implementation, and more importantly, the performance of video generation can hardly be preserved. In contrast, token merging [7] reduces the token size processed in attentions, decreasing computational complexity without requiring model re-training or hardware-specific adaptations.
@@
In STDiT [83], the attention bias influenced by input tokens in cross-attention layers leads to higher computational demands than in self-attention layers. As a result, developing an effective token merging method for cross-attention is crucial in video generation. However, existing token merging [6, 7, 34] applied to the cross-attention have shown suboptimal performance, and applying them to the self-attention in STDiT has observed video quality drops [6, 7].
5.2 Temporal Dimension Token Merging
@@
Figure 5: In attention layers of STDiT [83], two consecutive tokens are merged along the temporal dimension and subsequently unmerged after processing, reducing the token size by half and the computational complexity up to a quarter.
@@
Fig. 5 illustrates Temporal Dimension Token Merging. Based on the hypothesis that successive video frames exhibit similar values, two consecutive frames are merged over the temporal dimension by averaging, creating a single token without the overhead of calculating frame similarity. This reduces the size of tokens by half while preserving the essential temporal information. Consequently, it decreases the computation of self-attention by a factor of four, according to the self-attention’s quadratic complexity, $\mathcal{O}(n^{2})$. Similarly, it reduces the computation of cross-attention by half, based on its linear complexity, $\mathcal{O}(nm)$. Then, the output token processed through attentions replicates the dimensions for each frame, restoring them to their original size.
@@
Given the token $\boldsymbol{T}_{in}$ as an input for a self- or cross-attention with the dimension [$B$, $ST$, $C$], where $B$ denotes the batch size, $S$ is the number of pixel patches, $T$ is the number of frames, and $C$ is the channel dimension, the input token $\boldsymbol{T}_{in}$ is merged into $\boldsymbol{T}_{merged}$, with the index $i$, as:
$\boldsymbol{T}_{merged}=TDTM_{merge}(\boldsymbol{T}_{in})$ (6)
$TDTM_{merge}(\boldsymbol{T})[i]=\frac{1}{2}(\boldsymbol{T}[:,i,:]+\boldsymbol{T}[:,i+1,:])$ (7)
@@
From this, two adjacent tokens are merged along the temporal dimension, producing the merged token $\boldsymbol{T}_{merged}$ of [$B$, $ST/2$, $C$], which is processed by self- or cross-attention.
@@
$\boldsymbol{T}_{unmerged}=TDTM_{unmerge}(\text{Attention}(\boldsymbol{T}_{merged}))$ (8)
$TDTM_{unmerge}(\boldsymbol{T})[2i]=\boldsymbol{T}[:,i,:]$ (9)
@@
Figure 6: Example videos generated by On-device Sora and Open-Sora [83] (68 frames, 256×256).